OpenStudy (anonymous):

HELP ASAP, WILL GIVE MEDAL!! Part 1: Explain, in complete sentences, whether or not the expression x2 – 121 fits one of the special patterns. If so, which one? (2 points) Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

OpenStudy (anonymous):

Part 1: the pattern is a^2 - b^2 = (a-b)(a+b) I don't understand what parts 2 & 3 ask.

OpenStudy (anonymous):

part 1... 121 can be written as (11)^2 which allows us to write x^2 - (11)^2. This form can be simplified to a^2 - b^2 which can be written as (a+b)(a-b) part 2... since \[x^{2} - 121 = 0\]we need two number that add up to 0 and multiply to give -121 part 3... we can write x^2-121 as \[x^{2} - 121 = 0\]\[x^{2}=121\]\[x=\pm11\]therefore the factors are (x+11)(x-11) since 11-11=0 and 11 x -11 = -121